OpenStudy (anonymous):

3 - a =12 solve for a 4

OpenStudy (anonymous):

\[\frac{3}{4}a=12\]

OpenStudy (anonymous):

The first thing you should do is multiply whatever is on the right side by the denominator on the left.

OpenStudy (anonymous):

that would give you \(3a=48\)

OpenStudy (anonymous):

That leaves you with 3a=12(4) 12 times 4 equals 48 3a=48 48/3 = 16 a=16